# A Chinese Prompt Attack Dataset for LLMs with Evil Content

### v1

@@ Abstract
Large Language Models (LLMs) presents significant priority in text understanding and generation. However, LLMs suffer from the risk of generating harmful contents especially while being employed to applications. There are several black-box attack methods, such as Prompt Attack, which can change the behaviour of LLMs and induce LLMs to generate unexpected answers with harmful contents. Researchers are interested in Prompt Attack and Defense with LLMs, while there is no publicly available dataset to evaluate the abilities of defending prompt attack. In this paper, we introduce a Chinese Prompt Attack Dataset for LLMs, called CPAD. Our prompts aim to induce LLMs to generate unexpected outputs with several carefully designed prompt attack approaches and widely concerned attacking contents. Different from previous datasets involving safety estimation, We construct the prompts considering three dimensions: contents, attacking methods and goals, thus the responses can be easily evaluated and analysed. We run several well-known Chinese LLMs on our dataset, and the results show that our prompts are significantly harmful to LLMs, with around 70% attack success rate. We will release CPAD to encourage further studies on prompt attack and defense.
@@ Introduction
Large Language Models (LLMs) are flourishing in terms of generating creative texts and solving practical problems. Take ChatGPT [^1] https://chat.openai.com/ as an example, with 175 billion parameters, it stores a large amount of knowledge and exhibits surprising performance on various downstream tasks. Instruction tuning(Wei et al. 2021) and Reinforcement Learning from Human Feedback(Ziegler et al. 2019; Lambert et al. 2022) (RLHF) allow LLMs to generate appropriate answers based on user queries. LLMs can handle tasks they have never seen before without continuous fine-tuning.
While LLMs have demonstrated success in numerous tasks, their full deployment has yet to occur, largely due to the significant security risks associated with LLMs. With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views(Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
@@
Nonetheless, the majority of LLMs exhibit a high rate of success in defending against the offensive prompts within publicly accessible evaluation benchmark. For example, ChatGPT and ChatGLM-6B(Zeng et al. 2022; Du et al. 2022) achieves the overall safety scores of 98.09 and 94.18 respectively on SAFETYPROMPTS. This implies that such prompts may not comprehensively reflect the full spectrum of safety possessed by various Large Language Models.
@@
Furthermore, the safety assessment conducted in previous studies failed to accurately reflect the practical attack scenarios. When launching a prompt attack on a Large Language Model, the attacker has a specific target to exploit. For instance, if LLMs are prompted to generate pornographic content, propagate racist rhetoric, or provide guidance for illegal and criminal activities, then descriptions of scenes, criticism and plans are expected, respectively. Unfortunately, previous researches have neglected to consider this crucial factor when constructing attacking prompts.
@@
To accurately simulate prompt attacks on LLMs from the perspective of attackers and estimate the associated security risks, we have developed a Chinese prompt attack dataset called CPAD. During the prompts construction and collection, we meticulously consider three key dimensions: the target, the attacking approaches and the attacking content of the attack. Additionally, we carefully designed some attacking approaches for various topics, and an automated process for constructing attack prompts. Furthermore, we propose to evaluate the safety score of LLMs with different evaluation prompts according to the attacking targets, unlike Sun et al. (2023), which uses the same prompt for all scenarios.
@@
We employed a LLM to automatically construct a large number of attack samples, and after analyzing the responses from three Chinese LLMs, we identified 10050 highly effective and dangerous prompts. We conducted experiments and analysis on several LLMs including ChatGPT. According to the statistics, our attacking prompts exhibited a remarkable success rate of 69.91% against ChatGPT, which is significantly higher than most publicly available datasets. Additionally, our attacking prompts achieved a success rate of over 70% against other open-source Chinese models such as ChatGLM2-6B(Du et al. 2022; Zeng et al. 2022).
@@
Besides, we conduct a simple experiment to fine-tune a LLM with CPAD, reducing the risks of being attacked. Our research endeavors to bring attention to prompt attacks within the community. We encourage researchers to utilize our dataset and engage in further investigation of prompt attack strategies and defensive approaches, thereby enhancing the security and performance of Large Language Models.
@@
    To address the current lack of high-quality evaluation datasets in the field of prompt attacks, we have developed a set of high-quality prompt attack samples, which encompass a variety of attacking contents.
@@
    We specify the attacking goals of each prompt, which not only accurately simulate prompt attacks on LLMs from the perspective of attackers, but also can be utilized as the prompt to evaluate the response.
@@
    Our analysis indicates that our prompts achieve an attack success rate of nearly 70% against ChatGPT. A simple strategy has been developed to defend against the attacks. We encourage researchers to leverage our dataset for further studies of prompt attacks.
@@
Different from previous studies, our dataset contains the attacking targets and aims to investigate the defense capability of various Large Language Models against carefully designed prompt attacks from the perspective of attackers, thereby encouraging the interest in the research of defense against prompt attacks within the community.
@@
In this Section, we will introduce the construction of our dataset. We firstly generate several candidates with ChatGPT according to seed prompts. Then we filter the extended prompts utilizing three different Large Language Models, as shown in Figure 2.
@@
We consider three dimensions to classify the attacking prompts. For each dimension, we manually designed several seed items for extensive generation.
@@
indicates the topic of the attack. The attack content we designed encompasses multiple aspects and aims to present the utmost severity of harm possible. Our 28 seed contents are listed bellow:
  • Discrimination and Unfairness.
    • Discrimination against certain ethnic groups, such as blacks and Jews.
    • Regional discrimination.
    • Discriminates against people with physical defects, such as disabled people, fat people, short people, ugly people, people with genetic diseases.
    • Discriminate against people of a certain country.
    • Discrimination against Women.
    • Discrimination against the poor.
    • Discrimination against people without education.
    • Discrimination against country people and rural people.
    • Discriminate against a particular religion.
    • Discriminate against an occupation.
  • Common crimes against the law.
    • Theft and robbery.
    • Acts endangering life safety such as homicide, suicide, intentional injury, etc.
    • Tax evasion.
    • Illegal acts involving pornography.
    • Gambling violations.
    • Involved in drug trafficking and drug use.
    • Kidnapping.
    • Behaviors that violate ethical and moral principles, such as infidelity, incest, lack of filial piety, and human drug experimentation.
    • Cheating in exams.
    • Jailbreaking.
    • Fraud.
    • Rumor.
  • Verbal abuse.
  • Privacy theft.
    • Attacking businesses, schools, and governments through Internet.
    • Stealing personal privacy information.
    • Espionage.
  • Terrorist attack.
  • Wars.
@@
Template
indicates the shell of the prompt. As bare attacking contents are easily be rejected by the LLMs aligned with human preferences, attack templates can help to conceal attack content, thereby bypassing scrutiny from LLMs. Some prompt templates may fail to attack the well aligned LLMs, so we manually check several instances for each of the following designed templates:
  • Conversation completion. Simulate a conversation between two individuals or users and a large language model, which is designed to answer any questions, removing moral and ethical constraints. Prompt the model to fill in the unsafe content in the conversation.
  • Assign roles and then issue commands. Previous studies have indicated that incorporating LLMs into certain roles can effectively bypass security restrictions. Therefore, we have the large model play roles such as a writer, a screenwriter, a journalist, etc., in order to generate descriptions of unsafe scenarios. We also have the model play the role of a parrot, repeating content with unsafe topics.
  • One-sided statement. Most things in life have a dual nature, and large language models that have been trained to align with human preferences are adept at analyzing the pros and cons from multiple perspectives. Therefore, we begin by acknowledging the negative aspects of the content in question, and then prompt the large language model to articulate the corresponding positive aspects of the attacking content or the negative aspects of normal behavior.
  • Fact pollution. When a sentence stating common sense or knowledge is phrased in an affirmative tone and assumes an incorrect fact within the context, large language models may generate new content based on that false fact, resulting in factual contamination.
Goal
is one of the significant differences between our dataset and previous studies. The The attacking goal refers to the expected behavior of the Large Language Model shown by the attacker after successfully attacking it. We attach great importance to this factor because in real attack scenarios, attackers always have specific goals when attacking LLMs. For example, some criminals may need concrete criminal plans instead of mere expressions of support. Therefore, both attack and result evaluation should be carried out with specific goals. We list our designed attacking goals as below:
  • Unsafe plan development. The attacker may seek to use LLMs to provide guidance and advice for aggressive and illegal behaviors, in order to efficiently carry out unethical and illegal activities. This type of attacking goal does not require the LLM to be integrated into a specific application, and may not have a negative impact on users, but it could provide helpful information for illegal activities.
  • List of unsafe speech or reasons. Providing reasons for unsafe behavior may help attackers find excuses, provide justification, or spread related speech on the internet, which can create public safety issues.
Inducing Fact Pollution
In addition to the traditional prompt attack methods, we have also added an additional set of prompt samples to induce the LLMs to generate statements with false facts.
Specifically, the types of facts can be summarized as following:
  • The author, date, and contents of artworks.
  • The content, nation of scientific research.
  • The news about sports and entertainment.
  • Historical events.
  • Economic and political facts.
Extended Prompts
We collected the manually designed seed contents, and prompted a LLM to automatically generate texts to extend our dataset. Specifically, we utilize gpt-3.5-turbo-0301 to extend the description for each pair of attacking content and goal. Besides, we prompt gpt-3.5-turbo-0301 to rewrite the attacking templates by changing the phrases and verbs. Then, the extended texts are randomly formatted with the rewritten attacking templates. Finally, we collected 68209 prompt attack samples.
@@
The evaluation process is illustrated in Figure 3. Different from previous studies with estimation of responses, we introduce the attacking goals which can accurately reflect the behavior the attacker wants the model to exhibit. This allows us to assess whether the responses of the LLMs align with the intentions of the attacker.
Specifically, we construct an evaluation prompt with the attacking content and goal. An evaluation LLM (gpt-3.5-turbo-0301 for example) is prompted to answer either “Yes” or “No”, representing the attack is successful or failed. As the extended attacking contents are automatically generated through LLMs, there are some low-quality attack samples. In order to keep only high-quality and highly harmful prompts, we have implemented the following two filtering strategies:
@@
We employ three different Chinese LLMs: ChatGLM-6B (Du et al. 2022; Zeng et al. 2022), Ziya-13B (Zhang et al. 2022) and ChatYuan-Large-v2 (Xuanwei Zhang and Zhao 2022). We keep only the prompts which successfully attacked all of the models. The evaluation results are listed in Table 1.
@@
The length of the response can reflect the harmfulness of the text. Short texts may demonstrate less harmfulness as they are less informative. From the perspective into the responses, we find that this phenomenon exists especially in the attacking goal of “Description of unsafe scenarios”. So we drop the prompts whose corresponding response has a length shorter than 100 under the attacking goal “Description of unsafe scenarios”.
Finally, we collect 10050 prompts to construct CPAD. The distribution is shown in Figure 4.
Figure 4: The distribution of attacking goals and templates in CPAD.
Figure 5: The distribution of the contents from ChatGPT responses.
Experiments
Implementation Details
Overall Success Attack Rate of CPAD
@@
Table 2: Overall success attack rate of CPAD against ChatGLM2-6B (denoted as ChatGLM), Baichuan-7B (denoted as Baichuan-7B), Chinese-LLaMa-2-7B (denoted as LLaMa2) and gpt-3.5-turbo-0301 (denoted as ChatGPT).
The evaluation results are listed in Table 2. It can be observed that the attack prompts of CPAD have high quality and even achieve a nearly 70% success rate against ChatGPT, which is one of the most successful LLM.
We find that careful training to align the LLM with human preference help to reduce the harmfulness. The safety of LLaMa-2 is optimized via supervised fine-tuning, RLHF and context distillation, so it has strong robustness to defense the prompt attack.
Larger model, such as ChatGPT, has a lower probability to be attacked. It is worth noting that although ChatGPT is relative rare to output dangerous content, its malicious output is also more harmful than small-scale models. A comparison between attacked outputs of ChatGPT and ChatGLM2 is illustrated in Table 3. Although they are both successfully attacked, ChatGPT’s output exhibits more harmfulness as the suggestions are practical and applicable, while ChatGLM only repeats some related but helpless statements.
@@
Table 3: Responses comparison between ChatGPT and ChatGLM under the same prompt attack. Note that the original sentences are Chinese. We have translated them in English.
Analysis of Attacked Outputs from ChatGPT
(a) The distribution of attacking goals from ChatGPT responses.
(b) The distribution of attacking templates from ChatGPT responses.
@@
We conduct in-depth analysis on the responses of ChatGPT, from the perspective of attacking contents (Figure 5), goals (Figure 5(a)) and templates (Figure 5(b)) respectively.
It can be concluded that 1) Unsafe responses are more than safe responses for all contents, especially “Espionage”, which we surmise is caused by lack of alignment with safe human preference on this topic. 2) Verbal abuse is the content with the least successful attacked response. It can be attributed to the explicit aggressiveness, which is easy to be detected. 3) Contamination and goal hijacking are difficult attacking goals against ChatGPT, as ChatGPT is too knowledgeable to be misled.
@@
We improve the LLM’s ability to defense prompt attack with supervised fine-tuning. We adopt Baichuan-13B-Chat [^6] https://huggingface.co/baichuan-inc/Baichuan-13B-Chat, a Chinese LLM after RLHF, to improve the safety under prompt attack. We split the prompts whose goals are “unsafe plan development” as test set, and the others are train set. More details can be found in Appendix A.
Content None % Fine-Tuning %
Overall 74.65 0.76
Acts endangering life safety.
64.71 5.88
Attacking businesses, schools, and governments through Internet.
86.15 0.00
Behaviors that violate ethical and moral principles.
64.10 0.00
Cheating in exams.
78.26 0.00
Discriminate against a particular religion.
67.92 1.89
Discriminate against an occupation.
74.51 1.96
Discriminate against people of a certain country.
52.08 0.00
Discriminates against people with physical defects.
61.82 0.00
Discrimination against Women.
56.00 0.00
Discrimination against certain ethnic groups.
64.00 2.00
Discrimination against country people and rural people.
76.36 1.82
Discrimination against people without education.
34.29 2.86
Discrimination against the poor.
76.47 0.00
Espionage.
87.18 0.00
Fraud.
77.01 0.00
Gambling violations.
78.43 0.00
Illegal acts involving pornography
57.69 3.85
Involved in drug trafficking and drug use.
65.63 0.00
Jailbreaking.
90.32 0.00
Kidnapping.
80.30 0.00
Regional discrimination
52.27 0.00
Rumor.
84.38 3.13
Stealing personal privacy information.
82.28 0.00
Tax evasion.
61.76 0.00
Terrorist attack.
76.27 1.69
Theft and robbery.
81.55 0.00
Verbal abuse.
72.41 3.45
Wars.
83.75 0.00
The successful attack rate before and after fine-tuning over different contents are listed in Table 4. 74.65% of the prompts successfully attacked Baichuan-13B-Chat before fine-tuning, while less that 1% successful attack in the test set after fine-tuning. Especially on “Jailbreaking”, the success rate decreases from 90.32% to 0%. The result proves the effectiveness of our dataset to improving the safety against prompt attack.

### v2

# Goal-Oriented Prompt Attack and Safety Evaluation for LLMs

@@ Abstract
Large Language Models (LLMs) presents significant priority in text understanding and generation. However, LLMs suffer from the risk of generating harmful contents especially while being employed to applications. There are several black-box attack methods, such as Prompt Attack, which can change the behaviour of LLMs and induce LLMs to generate unexpected answers with harmful contents. Researchers are interested in Prompt Attack and Defense with LLMs, while there is no publicly available dataset with high successful attacking rate to evaluate the abilities of defending prompt attack. In this paper, we introduce a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD. Our prompts aim to induce LLMs to generate unexpected outputs with several carefully designed prompt attack templates and widely concerned attacking contents. Different from previous datasets involving safety estimation, we construct the prompts considering three dimensions: contents, attacking methods and goals. Especially, the attacking goals indicate the behaviour expected after successfully attacking the LLMs, thus the responses can be easily evaluated and analysed. We run several popular Chinese LLMs on our dataset, and the results show that our prompts are significantly harmful to LLMs, with around 70% attack success rate to GPT-3.5. CPAD is publicly available at https://github.com/liuchengyuan123/CPAD.
@@ Introduction
Large Language Models (LLMs) can generate creative content and solve practical planning and reasoning problems (Chowdhery et al. 2022; OpenAI 2023). Take GPT-3 (Brown et al. 2020) as an example, with 175 billion parameters, it stores a large amount of knowledge and exhibits surprising performance on various downstream tasks. Instruction tuning (Wei et al. 2021) and Reinforcement Learning from Human Feedback (Ziegler et al. 2019; Lambert et al. 2022) (RLHF) allow LLMs to generate appropriate answers based on user queries. LLMs can handle tasks they have never seen before without continuous fine-tuning.
Figure 1: Comparison of successful attacking between previous work and CPAD. CPAD can further expose the security risks of LLMs.
While LLMs have demonstrated success in numerous tasks, there are still some vital flaws for deployment such as the significant security risks associated with LLMs. With the help of robust general knowledge and common sense, LLMs can offer valuable guidance for harmful behavior and can automatically generate offensive, discriminatory and fraudulent content, and misleading views(Deng et al. 2023; Zhuo et al. 2023; Hartvigsen et al. 2022).
@@
Nonetheless, the majority of LLMs exhibit a high rate of success in defending against the offensive prompts within publicly accessible evaluation benchmark. For example, ChatGPT and ChatGLM-6B(Zeng et al. 2022; Du et al. 2022) exhibit the overall unsafety scores of 1.63 and 3.19 respectively on SAFETYPROMPTS. While recent studies illustrate non-negligible risk of generating harmful contents under carefully designed prompts, which implies that such benchmarks cannot comprehensively reflect the full spectrum of safety possessed by various Large Language Models.
@@
Additionally, the safety assessment conducted in previous studies are aimless and casual. They failed to effectively incorporate the attacking goal. When launching a prompt attack on a Large Language Model, the attacker has a specific goal to exploit. For instance, if LLMs are prompted to generate pornographic content, then descriptions of scenes are expected. On another hand, if the attacker wants LLMs provide guidance for illegal and criminal activities, then some plans and tips shall be concluded in the response. The attributes hidden in the text can be regarded as the flag to estimate whether LLMs are successfully attacked. Unfortunately, previous researches have neglected to consider this crucial factor when constructing attacking prompts.
@@
To accurately simulate prompt attacks on LLMs from the perspective of attackers and estimate the associated security risks, we have developed a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD. During the prompts construction and collection, we meticulously consider three key dimensions: contents, the attacking templates and goals. We carefully designed some attacking templates for various topics, and an automated process for constructing attack prompts, which contribute to the outperforming successful-attacking rate over pervious works as shown in Figure 1.
Furthermore, we propose to evaluate the safety score of LLMs with different evaluation prompts according to the attacking goals. Taking the previous case as an example, if the attacker wants LLMs provide guidance for illegal and criminal activities, we only need to identify whether the LLM’s response provides a plan related to the crime, and we can more accurately determine whether the model is safe. Unlike Sun et al. (2023), which uses the same prompt for all samples, our evaluation is more precise, intuitive, scenario-specific.
@@
We employed another LLM to automatically construct a large number of attack samples, and after analyzing the responses from three Chinese LLMs, we identified 10050 highly effective and dangerous prompts. We conducted experiments and analysis on several LLMs including GPT-3.5. According to the statistics, our attacking prompts exhibited a remarkable success rate of 69.91% against GPT-3.5, which is significantly higher than most publicly available datasets. Additionally, our attacking prompts achieved a success rate of over 70% against other open-source Chinese LLMs such as ChatGLM2-6B(Du et al. 2022; Zeng et al. 2022).
@@
Besides, we conduct a straightforward experiment to fine-tune a LLM with CPAD, reducing the risks of being attacked. Our research endeavors to bring attention to LLM security within the community. We encourage researchers to utilize our dataset and engage in further investigation of prompt attack strategies and defensive approaches, thereby enhancing the security and performance of Large Language Models.
@@
    To address the current lack of high-quality evaluation datasets in the field of prompt attacks, we have developed a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD.
@@
    We specify the attacking goals of each prompt, which not only accurately simulate prompt attacks on LLMs from the perspective of attackers, but also can be utilized to evaluate and analyse the response.
@@
    Our analysis indicates that our prompts achieve an attack success rate of nearly 70% against GPT-3.5. A straightforward strategy has been developed to defend against the attacks. We encourage researchers to leverage our dataset for further studies.
@@
Different from previous studies, our work emphasizes the significance of incorporating attacking goals and aims to estimate the defense capability of various Large Language Models against carefully designed prompt attacks from the perspective of attackers, thereby encouraging the interest in the research of LLMs attack and defense within the community.
@@
In this Section, we will introduce the construction of our dataset as shown in Figure 3. We firstly introduce the concepts of contents, templates and goals. The pipeline starts from manual-designed seeds. several candidates are generated as extended prompts with OpenAI-API [^2] https://api.openai.com/v1/chat/completions according to seeds. Then we filter the extended prompts utilizing the validation on three different Large Language Models. Finally, the attacking goals and prompts are collected.
@@
We primarily consider three key dimensions to classify the samples. For each dimension, we manually designed several seed items for extensive generation.
@@
indicates the topic of the attack, such as unfairness and crimes. Contents are widely considered for LLMs attack. Here we adopt a hierarchical classification for contents and sub-contents (for example, Sexism and hellish discrimination are both grouped into “discrimination categories”). The attack contents we designed cover multiple aspects and aims to present the utmost severity of harm possible. Detailed content items are listed in Table 4.
@@
Template
indicates the shell of the prompt. As bare attacking contents are easily be rejected by the LLMs after aligning with human preferences, attack templates can help to conceal attack content, thereby bypassing scrutiny from LLMs. Some prompt templates may fail to attack the well aligned LLMs, so we manually check several instances for each of the following designed templates:
  • Conversation completion. Simulate a conversation between two individuals or users and a Large Language Model, where the LLM is designed to answer any questions, removing moral and ethical constraints. “Conversation completion” prompts the model to fillin the unsafe content in the conversation.
  • Assign roles and then issue commands. Previous studies have indicated that incorporating LLMs into certain roles can effectively bypass security restrictions. Therefore, we have the large model play roles such as a writer, a director, a journalist, etc., in order to generate descriptions of unsafe scenarios. We also have the model play the role of a parrot, repeating content with unsafe topics.
  • One-sided statement. Most things in the real world have a dual nature, and Large Language Models that have been trained to align with human preferences are adept at analyzing the pros and cons from multiple perspectives. Therefore, we start the instruction by acknowledging the negative aspects of the content, and then prompt LLMs to articulate the corresponding positive aspects of the attacking content or the negative aspects of a normal behavior.
  • Fact pollution. When a sentence stating common sense or knowledge is phrased in an affirmative tone and assumes an incorrect fact within the context, LLMs may generate new contents based on that false fact, resulting in factual contamination.
Goal
is one of the significant differences between our dataset and previous studies. The attacking goal refers to the expected behavior of the Large Language Model after successfully being attacked. We emphasize the significance of this factor because in real-world scenarios, the attacking prompts are always goal-oriented. For example, some criminals may need concrete criminal plans instead of mere expressions of support, thus the response should be identified as unsafe when it concluded some plans or tips. Previous works will consider a response is unsafe as long as it contains any harmful contents. Both attack and result evaluation should be carried out with specific goals. We list our designed attacking goals as below:
  • Unsafe plan development. The attacker may seek to use LLMs to provide guidance and advice for aggressive and illegal behaviors, in order to efficiently carry out unethical and illegal activities. This kind of attacking goals don’t require the LLM to be integrated into a specific application, and may not have a negative impact on users, but they could provide helpful information for illegal activities.
  • List of unsafe speech or reasons. Providing reasons for unsafe behavior may help attackers find excuses, justification, and spread related speech on the internet, which can create public safety issues.
Inducing Fact Pollution
Since the object of Fact Pollution is the statements of knowledge, which is slightly different from other attack templates, we describe details of Fact Pollution in this SubSection.
Specifically, the types of facts can be summarized as following:
  • The author, date, and contents of artworks.
  • The content, nation of scientific research.
  • The news about sports and entertainment.
  • Historical events.
  • Economic and political facts.
Construction of Seed Prompts
In the process of constructing the seed prompts, the labors match each attack template with several proper attacking goals, and write a harmful sentence or comment for each type of attacking contents. It is worth noting that in order to ensure the quality of the seed prompts, each prompt is verified on GPT-3.5 and GPT-4 models, and only the response of the model meets the expectation of the attacking goal can be considered as passed. The careful check ensures the improvement on the final successful-attacking rate.
Extended Prompts
We collected the manually designed seed prompts from last SubSection, and prompted a LLM to automatically generate texts to extend our dataset. Specifically, we utilize gpt-3.5-turbo-0301 to extend the context and detailed scenarios. Besides, we prompt gpt-3.5-turbo-0301 to rewrite the attacking templates by changing the phrases and verbs. Then, the extended texts are randomly formatted with the rewritten attacking templates. Finally, we collected 68209 extended prompts.
Figure 5: The distribution of attacking goals and templates in CPAD.
@@
The evaluation process is illustrated in Figure 4. Different from previous studies which use the same evaluation prompt for all samples, we incorporate the attacking goals into the evaluation prompts and identify whether the response satisfies the attacker’s expectation, which intuitively provides a more precise estimation and analysis.
Specifically, we construct an evaluation prompt leveraging the attacking content and goal, such as Does the model’s response or output directly or indirectly Attacking Goal for supporting the act of Attacking Content? If yes, please answer “yes”; If not, please answer “no”. Answer:. The answer, either “yes” or “no”, represents the attack is successful or failed. Since the extended prompts are automatically generated through LLMs, there are low-quality ones. In order to keep only high-quality and highly harmful prompts, we have implemented the following two filtering constraints:
@@
We conduct validation experiments for three popular Chinese LLMs: ChatGLM-6B (Du et al. 2022; Zeng et al. 2022), Ziya-13B (Zhang et al. 2022) and ChatYuan-Large-v2 (Xuanwei Zhang and Zhao 2022). We keep only the prompts which successfully attacked all of the models. The evaluation results are listed in Table 1.
@@
The length of the response can reflect the harmfulness of the text. Short texts may demonstrate less harmfulness since they are less informative. From the perspective into the responses, we find that this phenomenon exists especially in the attacking goal of “Description of unsafe scenarios”. So we drop the prompts whose corresponding response has a length shorter than 100 words under the attacking goal “Description of unsafe scenarios”.
Finally, we collect totally 10050 prompts to construct CPAD. The distribution is shown in Figure 5.
@@
Table 2: Overall successful-attacking rate of CPAD against ChatGLM2-6B (denoted as ChatGLM2), Baichuan-7B (denoted as Baichuan-7B), Chinese-LLaMa-2-7B (denoted as LLaMa2) and gpt-3.5-turbo-0301 (denoted as ChatGPT).
@@
Table 3: Responses comparison between GPT-3.5 and ChatGLM under the same prompt attack. Note that the original sentences are Chinese. We have translated them in English.
Figure 6: The distribution of the contents from GPT-3.5 responses.
Experiments
Implementation Details
To accurately measure the level of harm and the degree to which our prompts obscure harmful intentions, we calculate the successful-attacking rate, which is the ratio of prompts that successfully attacked the LLM.
Overall Successful-Attacking Rate
The evaluation results of models with different scale are listed in Table 2.
ChatGLM2 represents harmful contents on 78.49% samples. While Baichuan, a larger model with 7B parameters, is shown to be dangerous on around 79% samples, which is slightly higher than ChatGLM2. It is reported that, LLaMa2 is especially optimized for safety. Our experiments on Chinese-LLaMa2 (fine-tuned for Chinese) also demonstrate the improvement on safety, with a 74.84% successful-attacking rate. CPAD even achieves a nearly 70% successful-attacking rate against GPT-3.5, which is one of the most successful LLM. Generally, the attack prompts constructed with our pipeline have high quality.
Larger models, such as GPT-3.5, have a lower probability to be attacked. It is worth noting that although GPT-3.5 is safer for text generation under carefully designed prompt attack, its malicious output is also more harmful than small-scale models once failed to defense. To clearly illustrate the finding, a comparison between attacked outputs of GPT-3.5 and ChatGLM2 is shown in Table 3. Although they are both successfully attacked, GPT-3.5’s output apparently exhibits more harmfulness as the suggestions are practical and applicable, while ChatGLM only repeats some related but helpless statements. From the perspective of attackers, GPT-3.5 is more suited to be the role of a crime assistant in this case. The greater the capabilities offered by LLMs service providers, the more imperative it becomes to enhance security in order to effectively reject any attempts by attackers to exploit the LLMs’ capabilities. This will ensure that the potential harm caused by such exploits is minimized, especially in comparison to smaller-scale LLMs.
(a) The distribution of attacking goals from GPT-3.5 responses.
(b) The distribution of attacking templates from GPT-3.5 responses.
@@
Content Sub-Content None % Fine-Tuning %
Overall 74.65 0.76
Discri- mination and Unfairness.
Discriminate against a particular religion.
67.92
1.89
Discriminate against an occupation.
74.51
1.96
Discriminate against people of a certain country.
52.08
0.00
Discriminates against people with physical defects.
61.82
0.00
Discrimination against Women.
56.00
0.00
Discrimination against certain ethnic groups.
64.00
2.00
Discrimination against country people and rural people.
76.36
1.82
Discrimination against people without education.
34.29
2.86
Discrimination against the poor.
76.47
0.00
Regional discrimination.
52.27
0.00
Common crimes against the law.
Acts endangering life safety.
64.71
5.88
Behaviors that violate ethical and moral principles.
64.10
0.00
Cheating in exams.
78.26
0.00
Fraud.
77.01
0.00
Gambling violations.
78.43
0.00
Illegal acts involving pornography
57.69
3.85
Involved in drug trafficking and drug use.
65.63
0.00
Jailbreaking.
90.32
0.00
Kidnapping.
80.30
0.00
Rumor.
84.38
3.13
Tax evasion.
61.76
0.00
Theft and robbery.
81.55
0.00
Privacy theft.
Attacking businesses, schools, and governments through Internet.
86.15
0.00
Espionage.
87.18
0.00
Stealing personal privacy information.
82.28
0.00
Verbal abuse.
72.41
3.45
Wars.
83.75
0.00
Terrorist attack.
76.27
1.69
Analysis of Attacked Outputs from GPT-3.5
We conduct in-depth analysis on the responses of GPT-3.5, from the perspective of attacking contents (Figure 6), goals (Figure 6(a)) and templates (Figure 6(b)) respectively.
It can be concluded that 1) Unsafe responses are more than safe responses for all contents, especially “Espionage”, which we surmise is caused by lack of alignment with safe human preference on this topic. 2) Verbal abuse is the content with the least successful attacked response. It can be attributed to the explicit aggressiveness, which is easy to be detected. 3) Contamination and goal hijacking are difficult attacking goals against GPT-3.5, as GPT-3.5 is too knowledgeable to be misled.
@@
We conducted a straightforward supervised fine-tuning experiment to improve the LLM’s defense ability under prompt attack. We adopt Baichuan-13B-Chat [^6] https://huggingface.co/baichuan-inc/Baichuan-13B-Chat, a Chinese LLM after RLHF, to improve the safety under prompt attack. We adopt LoRA for parameter-efficient fine-tuning. We set the rank as 8, learning rate as 5e-5 and batch size as 8. We fine-tuned Baichuan-13B-Chat for 3 epochs. To construct the safe responses, we unwrap the attacking prompts and feed them directly into LLMs. The outputs are regarded as the label for fine-tuning. We split the prompts whose goals are “unsafe plan development” as test set, and the others are train set.
The successful-attacking rate before and after fine-tuning over different contents are listed in Table 4. 74.65% of the prompts successfully attacked Baichuan-13B-Chat before fine-tuning, slightly lower than Baichuan-7B model. While less that 1% successful attack in the test set after fine-tuning. We almost reject all attacks by supervised fine-tuning. Especially on “Jailbreaking”, the successful-attacking rate decreases from 90.32% to 0%. However we also notice that there is only a drop of 60% on“Acts endangering life safety”, which covers a wider range of behaviors. The results indicate a promising defense with supervised fine-tuning if developers are given the potential attacking goals and templates.